OpenStudy (anonymous):

Which of the following statements is not true regarding operations with exponents? a.To divide powers with the same base, subtract the exponents. b.To subtract powers with the same base, divide the exponents. c.To multiply powers with the same base, add the exponents. d.To raise a power to a power, multiply the exponents.

OpenStudy (anonymous):

Exponent rules 1. \[4^{4} \times 4^{5} = 4^{4+5} = 4^{9}\] 2. \[\frac{ 4^{5} }{ 4^{2} } = 4^{5-2}\] 3. \[(5^{3})^{5} = 5^{3\times5}\]

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

a little, yes, thank you

OpenStudy (anonymous):

good examples were given of the rules of exponents.... which one was not given?

OpenStudy (anonymous):

1. is multiplying 2. is dividing 3. Is raising a power by a power They area ll the rules, so can you see which one is not given?

OpenStudy (anonymous):

b, thank you